Find the area under the normal curve between a. z = -1.22 and z = 2.32 b. z =1.27 and z= 1.77, c. z =2.0 and z = -0.50. 


1
Expert's answer
2020-12-24T16:21:06-0500

a.

"P(-1.22<Z<2.32)=P(Z<2.32)-P(Z\leq-1.22)"

"\approx0.98982956-0.11123244\approx0.8786"

"0.8786"


b.

"P(1.27<Z<1.77)=P(Z<1.77)-P(Z\leq1.27)"

"\approx0.96163643-0.89795768\approx0.0637"

"0.0637"


c.

"P(-0.50<Z<2.0)=P(Z<2.0)-P(Z\leq-0.50)"

"\approx0.97724987-0.30853754\approx0.6687"

"0.6687"